OpenStudy (anonymous):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for two different years: Number of years 1 2 3 Option 1 (amount in dollars) 1100 1210 1331 Option 2 (amount in dollars) 1100 1200 1300 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Part C:

OpenStudy (anonymous):

Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1? Explain your answer, and show the investment value after 20 years for each option. (4 points)

OpenStudy (anonymous):

I know A Option 1 Exponential Option 2 Linear B f(x)=1000+(1.1)x

OpenStudy (danjs):

The exponential option general form is.. value = A * (r)^t

OpenStudy (danjs):

you are given 3 data points, (t , value) You can use those to determine A and r

OpenStudy (danjs):

oh, when t=0, value = 1000

OpenStudy (danjs):

using that initial point (0 , 1000) \[1000 = A*r^0 = A*1 = A\]

OpenStudy (danjs):

A is 1000, the equation becomes, \[Value = 1000*r^t\]

OpenStudy (danjs):

use any other point for value and t, to calculate the r

OpenStudy (anonymous):

thats the answer for c? @danj5